OpenStudy (anonymous):

Mac deposited $918.78 in a savings account that earns 2.1% simple interest. What is mac's account balance after 10 years?

OpenStudy (anonymous):

initial = 918.78 interest= .021 time = 10 years so equation is... initial*e^(time*interest)

OpenStudy (anonymous):

so it would be 918.78(10*0.021)?

OpenStudy (anonymous):

no, (10*.021) has to be to the power of e. and e is equal to 2.7818... or something, it should be on your graphing calculator.

OpenStudy (anonymous):

I don't have a graphing calculator. What would E be? I got the rest but I have no idea what e is.

OpenStudy (anonymous):

with out e I got 192.94.

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

can you help me with this?

OpenStudy (anonymous):

I think it would be 918.78(10*0.021)=192.94+918.78= 1,111.72

OpenStudy (anonymous):

e is 2.71828

OpenStudy (anonymous):

it should be 1133.48

OpenStudy (anonymous):

if you multiply 10 by 0.021 it equals 0.21

OpenStudy (anonymous):

yes and that is to the power of e.

OpenStudy (anonymous):

no power because your answer isn't one of my choices

OpenStudy (anonymous):

@Hero can you help me with this please?